OpenStudy (anonymous):

In Mrs. Brown's Geometry class there are 15 boys and 10 girls. The students are presenting projects to the class. If Mrs. Brown selects students at random for presentations, what is the probability that the first 2 students chosen are girls?

OpenStudy (anonymous):

P ( 1st girl) = 10/ 25 = 2/5 P ( 2nd girl) = 9/24 = 3/ 8 => P (both girls) = 3/ 20

OpenStudy (anonymous):

Alternate solution that is probably not needed for a problem this simple, but might be of use in a more complicated problem......\[P(g,g)=\frac{\left(\begin{matrix}10 \\ 2\end{matrix}\right)\left(\begin{matrix}15 \\ 0\end{matrix}\right)}{\left(\begin{matrix}25 \\ 2\end{matrix}\right)}=\frac{\frac{10*9*8!}{2!8!}}{\frac{25*24*23!}{2!23!}}=\frac{45}{300}=\frac{3}{20}\]It's always cool to know more than one way to work on these.....